# Improving Pareto Set Learning for Expensive Multi-objective Optimization via Stein Variational Hypernetworks

### v2

@@ Introduction
We sample $K$ preference vectors $\bm{r}_{i}$ and fit them to the Pareto set model $h(r|\theta)$ to generate solutions $\{\bm{x}_{r_{i}}\}_{i=1}^{K}$. For a problem with $\bm{m}$ objectives, the objective functions are represented as $\mathcal{F}_{i}(\bm{x}_{r_{i}})=[f_{1}(\bm{x}_{r_{i}}),...,f_{m}(\bm{x}_{r_{i}})]$, where $\mathcal{F}_{i}$ is considered as a particle. The set of particles $\mathbf{F}={\{\mathcal{F}_{i}}\}^{K}_{i=1}$ serves as the initialization set. We iteratively move the particles towards the Pareto front using the update rule:
@@
$\bm{k}(\mathcal{F}_{i},\mathcal{F}_{j})=\sum^{m}_{n=1}\bm{k}_{n}\left(f_{n}(\bm{x}_{r_{i}}),f_{n}(\bm{x}_{r_{j}})\right).id_{n}$ (15)
Here $\bm{k}_{n}(\cdot)$ is the kernel for $n$ th dimension, $id_{n}$ is 1 in case $n$ is the index of the function $f_{n}$ that achieves the maximum formula (9), 0 otherwise.
@@
Discussion Intuitively, in SVH-PSL executes direct operations via the kernel to modify the gradient. It is well established that gradients are highly sensitive and can easily experience issues such as gradient explosion or vanishing. Figure 2 illustrates that when a particle is significantly affected by the presence of other particles, this interaction can facilitate a more rapid convergence. However, it may also lead to instability, potentially diverting the particle from reaching the global optimum. The implementation of our local kernel facilitates precise adjustments aimed at minimizing the potential for instabilities.
@@
Baseline Methods. We compare SVH-PSL with the current state of the art MOBO PFL methods including TS-TCH (Paria, Kandasamy, and Póczos 2020), USeMO-EI (Belakaria et al. 2020), MOEA/D-EGO (Zhang et al. 2009), TSEMO (Bradford, Schweidtmann, and Lapkin 2018), DGEMO (Konakovic Lukovic, Tian, and Matusik 2020), qParEGO (Knowles 2006), qEHVI (Daulton, Balandat, and Bakshy 2020), qNEHVI (Daulton, Balandat, and Bakshy 2021), DA-PSL (Lu, Li, and Zhou 2024) and PSL-MOBO (Lin et al. 2022).
@@
The Trade-Off Between Exploration and Exploitation. In Formula 14, the second term acts as the repulsion component, with the parameter $\alpha$ serving as a trade-off between exploration and exploitation. As shown in Figure 5, $\alpha$ significantly impacts performance. When $\alpha$ is small (left), the variance is high, indicating that insufficient repulsion makes it difficult for the particles to escape pseudo-optimal points, leading to instability. Conversely, a larger $\alpha$ (right) encourages exploration, driving the particles toward the optimal solution.
High Dimensional. Figure 6 shows that SVH-PSL performs well in high-dimensional conditions. Although optimization problems become increasingly challenging as dimensionality rises, SVH-PSL maintains its effectiveness. However, it also exhibits higher variance.

### v3

@@ Introduction
We sample $K$ preference vectors $\bm{r}_{i}$ and fit them to the Pareto set model $h(r|\theta)$ to generate solutions $\{\bm{x}_{r_{i}}\}_{i=1}^{K}$. For a problem with $\bm{m}$ objectives, the objective functions are represented as $\mathcal{F}_{i}(\bm{x}_{r_{i}})=[\hat{\bm{f}_{1}}(\bm{x}_{r_{i}}),...,\hat{\bm{f}_{m}}(\bm{x}_{r_{i}})]$, where $\mathcal{F}_{i}$ is considered as a particle. The set of particles $\mathbf{F}={\{\mathcal{F}_{i}}\}^{K}_{i=1}$ serves as the initialization set. We iteratively move the particles towards the Pareto front using the update rule:
@@
$\bm{k}(\mathcal{F}_{i},\mathcal{F}_{j})=\sum^{m}_{n=1}\bm{k}_{n}\left(\hat{\bm{f}_{n}}(\bm{x}_{r_{i}}),\hat{\bm{f}_{n}}(\bm{x}_{r_{j}})\right).id_{n}$ (15)
Here $\bm{k}_{n}(\cdot)$ is the kernel for $n$ th dimension, $id_{n}$ is 1 in case $n$ is the index of the function $\hat{\bm{f}_{n}}$ that achieves the maximum formula (9), 0 otherwise.
@@
Discussion. Intuitively, SVH-PSL executes direct operations via the kernel to modify the gradient. It is well established that gradients are highly sensitive and can easily experience issues such as gradient explosion or vanishing. Figure 2 illustrates that when a particle is significantly affected by the presence of other particles, this interaction can facilitate a more rapid convergence. However, it may also lead to instability, potentially diverting the particle from reaching the global optimum. The implementation of our local kernel facilitates precise adjustments aimed at minimizing the potential for instabilities.
@@
Baseline Methods. We compare SVH-PSL with the current state-of-the-art MOBO PFL methods, including TS-TCH (Paria, Kandasamy, and Póczos 2020), USeMO-EI (Belakaria et al. 2020), MOEA/D-EGO (Zhang et al. 2009), TSEMO (Bradford, Schweidtmann, and Lapkin 2018), DGEMO (Konakovic Lukovic, Tian, and Matusik 2020), qParEGO (Knowles 2006), qEHVI (Daulton, Balandat, and Bakshy 2020), qNEHVI (Daulton, Balandat, and Bakshy 2021), DA-PSL (Lu, Li, and Zhou 2024) and PSL-MOBO (Lin et al. 2022).
@@
The Trade-Off Between Exploration and Exploitation. In Formula 14, the second term acts as the repulsion component, with the parameter $\alpha$ serving as a trade-off between exploration and exploitation. As shown in Figure 5, $\alpha$ significantly impacts performance. When $\alpha$ is small (left), the variance is high, indicating that insufficient repulsion makes it difficult for the particles to escape pseudo-optimal points, leading to instability. Conversely, a larger $\alpha$ (right) encourages exploration, driving the particles toward the optimal solution. High Dimensional. Figure 6 shows that SVH-PSL performs well in high-dimensional conditions. Although optimization problems become increasingly challenging as dimensionality rises, SVH-PSL maintains its effectiveness. However, it also exhibits higher variance.